What is the smallest positive integer that must be multiplied to 60 to get a perfect cube?
Answer by Alan3354(69443)

What is the smallest positive integer that must be multiplied to 60 to get a perfect cube?
-----------
60 = 2*2*3*5
PC = 2*2*2*3*3*3*5*5*5 = 27000
27000/60 = 450